# Semantics-Oriented Multitask Learning for DeepFake Detection: A Joint Embedding Approach

### v1

@@ V-B2 Training on FFSC
As shown in Table V, SJEDD performs significantly better than the competing detectors on the four unseen DeepFake datasets. This shows the effectiveness and generality of the proposed semantics-oriented joint embedding in enhancing DeepFake detection across various training datasets. Additionally, all detectors demonstrate strong performance on FF++, with RECCE achieving an impressive AUC of $95.56\%$. This can be attributed to the relatively small domain gap between FF++ and FFSC, as a portion of images in these datasets share the same video source (i.e., YouTube). Consequently, RECCE can easily “shortcut” by learning signal-level features through face reconstruction to improve DeepFake detection on FF++. In contrast, the detection accuracy on DFDC is comparatively lower, consistent with the observations in Table II.
@@ V-C1 Ablation from the Model Perspective
We first analyze two key components of SJEDD: the joint embedding in Eq. (4) and the fidelity loss in Eq. (IV-C). To assess their effectiveness, we construct two SJEDD degenerates: one implemented solely by the image encoder to produce the raw scores (i.e., $\bm{s}(\bm{x})=\bm{f}_{\bm{\phi}}(\bm{x})$) and trained with the cross-entropy loss, and the other implemented by joint embedding and trained with the cross-entropy loss. Table VI shows the cross-dataset evaluation results, in which all model variants are trained on FFSC.
@@
It is clear from the table that joint embedding achieves noticeably better performance than directly predicting the label hierarchy from the input face image. This highlights the effectiveness of joint embedding in multitask learning, where the model can automatically adjust its capacity for each task through end-to-end optimization. Additionally, joint embedding facilitates task relationship modeling as distances between text embeddings in the learned feature space. By further incorporating the loss of fidelity, SJEDD achieves the best results. As discussed previously, this could be due to the favorable mathematical properties of the fidelity loss compared to the cross-entropy loss [44]. Fig. 4 illustrates the training dynamics of these two loss functions, where we find that the cross-entropy loss exhibits pronounced fluctuations, whereas the fidelity loss curve is more stable. This provides additional evidence that the fidelity loss may guide the optimization towards better local optima.
Figure 4: Training dynamics of the fidelity loss versus the cross-entropy loss.
(a) FFSC: Age
(b) FFSC: Expression (Smile)
(c) FFSC: Expression (Surprise)
(d) FFSC: Gender
(e) FFSC: Identity
(f) FFSC: Pose
(g) FF++: DF
(h) FF++: F2F
(i) FF++: FS
(j) FF++: NT
(k) CDF
(l) DFDC
Figure 5: Illustration of the proposed SJEDD trained on FFSC in making predictions at the face attribute and region levels. For FFSC (as an intra-dataset test), fake attributes/regions are highlighted in bold. For FF++, CDF, and DFDC (as cross-dataset tests), only the inference results are provided for reference. Zoom in for improved visibility.
@@
where $\|\cdot\|_{2}$ denotes the $\ell_{2}$-norm and $\tau$ is the same temperature parameter as in Eq. (4), and 5) optimizing SJEDD with a fixed set of loss weightings $\{1.0,0.1,0.1\}$ in Eq. (5) to prioritize the primary task manually.
@@
w/o Bi-level Optimization 88.90 92.04 79.94 83.75 86.16
@@
Table VII presents the cross-dataset results with several useful findings. First, freezing both the image encoder and the text encoder (i.e., using the pre-trained CLIP for zero-shot DeepFake detection) proves ineffective in this task. Fine-tuning the image encoder significantly boosts the detection performance, which validates the importance of vision-language alignment. Joint fine-tuning of the image and text encoders offers additional performance gains.
Additionally, SJEDD is fairly robust to changes in the format of the text templates, which experiences only slight performance drops when using simplified keyword-based templates. Using normalized similarity for raw score calculation negatively impacts the detection performance, indicating that some discriminative information may be captured in the magnitude of the visual embedding. Finally, dynamically adjusting the loss weightings and prioritizing the primary task through bi-level optimization proves advantageous, which also removes the need for cumbersome hyperparameter tuning.
@@ V-C2 Ablation from the Data Perspective
To evaluate the effectiveness of the proposed semantics-oriented dataset expansion technique, we first isolate the roles of its two steps: data augmentation and data relabeling. We then compare our method with the widely used Face X-ray data augmentation [21] and the human-involved data annotation in FFSC [41]. Tables VIII and IX present the results, where models are trained on FF+ [18] (including any dataset expansion if applicable) by default unless stated otherwise.
TABLE VIII: AUC results of SJEDD trained using different variants of the proposed semantics-oriented dataset expansion technique applied to FF++. “w/o relabeling” corresponds to the binary classification paradigm
@@
As shown in Table VIII, data relabeling alone marginally improves SJEDD’s generalization to unseen datasets. This happens because, in FF++, the relabeling process simply merges DF [1] and FS [83] into one category, and F2F [84] and NT [4] into another, which does not entirely fulfill the potential of semantics-oriented categorization of face manipulations. As a result, SJEDD reduces to a manipulation-oriented detector with a risk of overfitting. Training with data augmentation consistently improves detection generalizability, aligning with the findings in [21, 24, 30]. The best performance is achieved by the complete SJEDD, in which both data augmentation and relabeling are activated. The improvements arise because our proposed semantics-oriented dataset expansion method enables a single manipulation to alter multiple attributes and meanwhile group multiple manipulations under the same attribute category. This encourages SJEDD to learn transferable features across different manipulations that affect similar face attributes, rather than relying solely on manipulation-specific cues.
(a) FF++: DF
(b) FF++: F2F
(c) FF++: FS
(d) FF++: NT
(e) CDF
(f) FSh
(g) DF-1.0
(h) DFDC
Figure 6: Visualization of SJEDD activation maps. In each subfigure, the forged face image is displayed on the left, with the corresponding attention heatmap on the right. Best viewed in color.
TABLE IX: AUC results of SJEDD variants trained using different data augmentation methods
@@
As shown in Table IX, the Face X-ray-augmented SJEDD performs worse than the proposed SJEDD. Furthermore, when tested on CDF [46], DF-1.0 [47], and DFDC [48], the proposed SJEDD achieves results comparable to the FFSC-trained SJEDD, which is exposed to a larger set of real-world sophisticated face manipulations and more comprehensive global face attributes. This underscores the effectiveness of our method in automatically expanding existing DeepFake datasets and promoting their generality. Additionally, models trained on FF++ do not generalize well to FFSC compared to models trained on FFSC and tested on FF++. This suggests that FFSC presents a more challenging test set.
@@ V-D Interpretability
The proposed SJEDD via vision-language correspondence enhances model interpretability by providing human-understandable explanations. Specifically, we leverage SJEDD trained on FFSC [41] to illustrate this interpretability, as shown in Fig. 5. SJEDD accurately highlights the modified face attributes and regions with high probabilities. For example, in Fig. 5 (b), the detector identifies the face as fake and meanwhile indicates a likely expression alteration, pinpointing the mouth and lip as manipulated regions. This allows us to combine the predefined text templates and form a textual explanation such as “The face image is fake because the global face attribute of expression is altered, through manipulation of the local lip and mouth regions.” Regarding the cross-dataset results on FF++, CDF, and DFDC, SJEDD can still successfully interpret the binary classification results by predicting the authenticity of individual face attributes and regions. For example, on FF++, our detector accurately identifies DF and FS as instances of identity manipulation and F2F and NT as examples of expression forgery.
We also present SJEDD activation maps by Grad-CAM [91] in Fig. 6, where warmer regions indicate more significant importance. As can be seen, SJEDD consistently focuses on the face regions to identify forgeries, aligning with our primary design goal of semantics-oriented DeepFake detection.
@@ VI Conclusion and Discussion
We have described a semantics-oriented joint embedding DeepFake detection method called SJEDD, along with a dataset expansion technique. The core of SJEDD lies in its use of joint embedding, which reformulates model parameter splitting in multitask learning to a model capacity allocation problem, allowing for end-to-end bi-level optimization. We replaced the cross-entropy loss with the fidelity loss in bi-level optimization, which additionally boosts the performance. SJEDD also renders some degree of model interpretation by providing human-understandable explanations. We hope that the development of SJEDD will inspire researchers in related fields to rethink DeepFake detection from a semantic perspective.
The present study addresses DeepFake detection purely from a technical standpoint. Future directions in this field are multifaceted, involving advancements in both technology and policy. From a technological perspective, further exploration of a multimodal approach, which combines audio, visual, and textual analysis, is recommended to enhance detection accuracy. Detectors should also be able to operate in real-time on edge devices to support live analysis of streaming content in large-scale media platforms. Another important avenue for advancement is to develop detectors capable of continual learning and adaptation to emerging DeepFake techniques, potentially involving human oversight to maintain detection accuracy and efficacy. Proactive measures such as using blockchain technology and digital watermarking to authenticate original content should also be considered. From a policy perspective, it is essential to establish international standards and frameworks for DeepFake detection. This would involve collaboration between governments, tech companies, and research institutions. Laws and regulations should also be enacted to address the creation and distribution of DeepFake content, including penalties for misuse and support for victims.

### v2

@@ V-B2 Training on FFSC
As shown in Table V, SJEDD performs much better than the competing detectors on the four unseen DeepFake datasets. This shows the effectiveness and generality of the proposed semantics-oriented joint embedding in enhancing DeepFake detection. Additionally, all detectors demonstrate strong performance on FF++, with RECCE achieving an impressive AUC of $95.56\%$. This is due to the relatively small domain gap between FF++ and FFSC, as a portion of images in these datasets share the same source (i.e., YouTube). Consequently, RECCE may easily “shortcut” through face reconstruction. In contrast, the detection accuracy on DFDC is comparatively lower, consistent with the observations in Table II.
@@ V-C1 Ablation from the Model Perspective
We first analyze two key components of SJEDD: the joint embedding in Eq. (4) and the fidelity loss in Eq. (IV-C). Specifically, we construct two SJEDD degenerates: one implemented solely by the image encoder to produce the raw scores (i.e., $\bm{s}(\bm{x})=\bm{f}_{\bm{\phi}}(\bm{x})$) and trained with the cross-entropy loss, and the other implemented by joint embedding and trained with the cross-entropy loss. Table VI shows the cross-dataset evaluation results, in which all model variants are trained on FFSC.
(a) Pretrained CLIP
(b) SJEDD
Figure 4: Illustration of semantic relationships among tasks before and after semantics-oriented multitask learning. The relationships are shown using a correlation matrix, in which each entry represents the cosine similarity between two task-specific textual embeddings. Zoom in for improved visibility.
@@
It is clear from the table that joint embedding achieves noticeably better performance than directly predicting the label hierarchy from the input face image. This highlights the importance of adjusting model capacity for each task through end-to-end optimization. Additionally, joint embedding facilitates task relationship modeling. As shown in Fig. 4, initially, the pretrained CLIP text encoder does not adequately capture semantic similarities among tasks, treating most tasks equally. However, after optimizing for joint embedding, the relationships between tasks are effectively learned. For example, the concept of “expression” becomes more closely associated with “lip” and “mouth.” By further incorporating the fidelity loss, SJEDD achieves the best results. Fig. 5 illustrates the training dynamics of the two loss functions, where we find that the cross-entropy loss exhibits pronounced fluctuations, whereas the fidelity loss curve is more stable.
Figure 5: Training dynamics induced by the cross-entropy and fidelity losses.
(a) FF++: Real
(b) CDF: Real
(c) DFDC: Real
(d) FFSC: Age
(e) FFSC: Expression (Smile)
(f) FFSC: Expression (Surprise)
(g) FFSC: Gender
(h) FFSC: Identity
(i) FFSC: Pose
(j) FF++: DF
(k) FF++: F2F
(l) FF++: FS
(m) FF++: NT
(n) CDF
(o) DFDC
Figure 6: Illustration of the proposed SJEDD trained on FFSC in making predictions at the face attribute and region levels. For FFSC (as an intra-dataset test), fake attributes/regions are highlighted in bold. For FF++, CDF, and DFDC (as cross-dataset tests), only prediction results are provided for reference. Zoom in for improved visibility.
@@
where $\|\cdot\|_{2}$ denotes the $\ell_{2}$-norm and $\tau$ is the same temperature parameter as in Eq. (4), 4) optimizing SJEDD with a fixed set of loss weightings $\{1.0,0.1,0.1\}$ in Eq. (5) to prioritize the primary task manually [24, 25, 33, 26, 36], 5) optimizing SJEDD with fixed equal loss weightings $\{1.0,1.0,1.0\}$ in Eq. (5), and 6) training SJEDD with dynamic weighting average [105] without prioritizing the primary task.
@@
w/ Loss Weightings of {1.0, 0.1, 0.1} 88.90 92.04 79.94 83.75 86.16
w/ Loss Weightings of {1.0, 1.0, 1.0} 88.32 92.93 82.27 81.48 86.25
w/ Dynamic Weighting Average [105] 89.02 93.38 82.06 82.16 86.66
@@
Table VII presents the cross-dataset results. First, freezing both the image and text encoders proves ineffective. Fine-tuning the image encoder significantly boosts detection performance, which validates the importance of vision-language alignment. Joint fine-tuning of the image and text encoders offers additional performance gains. Besides, SJEDD is fairly robust to changes in text template. Using normalized similarity for raw score calculation negatively impacts detection performance, indicating that some discriminative information may be captured in visual embedding magnitudes. Finally, dynamically adjusting the loss weightings and prioritizing the primary task through bi-level optimization proves advantageous, which also removes the need for cumbersome hyperparameter tuning.
(a) FF++: Real
(b) CDF: Real
(c) DFDC: Real
(d) FFSC: Real
(e) FF++: DF
(f) FF++: F2F
(g) FF++: FS
(h) FF++: NT
(i) CDF
(j) FSh
(k) DF-1.0
(l) DFDC
Figure 7: Visualization of Grad-CAM activation maps. Each subfigure shows a forged face (left), baseline heatmap (middle), and SJEDD heatmap (right). Best viewed in color.
Since SJEDD is built upon CLIP, we also compare it against recent CLIP-based DeepFake detectors, including CLIP [47], VLFFD [35], GM-DF [93], FCG [94], CLIPPing [92], and RepDFD [95]. Table VIII presents the cross-dataset results, where SJEDD consistently outperforms these CLIP-based detectors, further demonstrating the non-trivial adaptation of CLIP in SJEDD for DeepFake detection.
TABLE VIII: AUC results of the proposed SJEDD against recent CLIP-based DeepFake detectors in the cross-dataset setting. All results are copied directly from the original papers except for the zero-shot CLIP
Method Venue Backbone CDF DFDC Mean AUC
CLIP [47] ICML2021 ViT-L 77.70 74.20 75.95
VLFFD [35] ArXiv2023 ViT-L 84.80
GM-DF [93] ArXiv2024 ViT-B 83.20 77.20 80.20
FCG-L [94] ArXiv2024 ViT-L 92.30 81.20 86.75
FCG-B [94] ArXiv2024 ViT-B 83.20 75.60 79.40
CLIPPing [92] ICMR2024 ViT-L 71.90
RepDFD [95] AAAI2025 ViT-L 89.90 81.00 85.45
SJEDD (Ours) ViT-B 92.22 84.14 88.18
Furthermore, though SJEDD is trained on a GAN-based dataset (i.e., FF++ [20]), we assess its ability to detect diffusion-based face forgery using DiffusionFace [106] and DiFF [107]. As shown in Table IX, SJEDD achieves strong performance and surpasses other state-of-the-art detectors. These results further validate the effectiveness of our semantics-oriented approach in learning more generalizable features beyond manipulation-based artifacts [30, 26].
TABLE IX: AUC results on the two diffusion-based DeepFake datasets. $\dagger$” indicates that we exclude purely synthesized faces in these datasets, as they are beyond the scope of our current study
Method DiffusionFace [106] DiFF [107] Mean AUC
MADD [28] 74.72 87.35 81.04
FRDM [30] 73.23 55.22 64.23
RECCE [24] 78.64 82.00 80.32
CADDM [26] 62.94 60.20 61.57
SO-ViT-B [44] 90.55 83.59 87.07
SJEDD (Ours) 93.70 84.64 89.17
@@ V-C2 Ablation from the Data Perspective
To evaluate the effectiveness of our semantics-oriented dataset expansion technique, we first isolate the contributions of its two main steps: data augmentation and data relabeling. We then compare our method with the widely used Face X-ray data augmentation [23] and the human-in-the-loop data annotation in FFSC [44].
TABLE X: AUC results of SJEDD trained using different variants of our dataset expansion technique applied to FF++. “w/o relabeling” corresponds to binary classification
@@
As shown in Table X, data relabeling alone marginally improves SJEDD’s generalization to unseen datasets. This happens because the relabeling process simply merges DF [1] and FS [45] into one category, and F2F [46] and NT [4] into another in FF++, which does not fully fulfill the potential of semantics-oriented categorization of face manipulations. As a result, SJEDD essentially reduces to a manipulation-oriented detector with a risk of overfitting. Training with data augmentation consistently improves detection performance, aligning with the findings in [23, 25, 31]. The best performance is achieved when both data augmentation and relabeling are activated. The improvements arise because our proposed semantics-oriented dataset expansion method enables a single manipulation to alter multiple attributes, and meanwhile groups multiple manipulations under the same attribute category. This encourages SJEDD to learn transferable features across different manipulations that affect similar face attributes, rather than relying solely on manipulation-specific cues.
TABLE XI: AUC results of SJEDD variants trained using different data augmentation methods
@@
As shown in Table XI, the Face X-ray-trained SJEDD performs worse than the proposed SJEDD. Furthermore, our detector achieves results comparable to the FFSC-trained variant, which is exposed to a larger set of real-world sophisticated face manipulations and more comprehensive global face attributes. This underscores the effectiveness of our technique in automatically expanding existing DeepFake datasets.
@@ V-D Interpretability
The proposed SJEDD via vision-language correspondence enhances model interpretability by providing human-understandable explanations. Specifically, we leverage SJEDD trained on FFSC [44] to illustrate this, as shown in Fig. 6. SJEDD accurately highlights the modified face attributes and regions with high probabilities. For example, in Fig. 6(e), the detector identifies the face as fake and meanwhile indicates a likely expression alteration, pinpointing the mouth and lip as manipulated regions. This allows us to combine the predefined text templates and form a textual explanation such as “The face image is fake because the global face attribute of expression is altered, through manipulation of the local lip and mouth regions.” As for images in FF++, CDF, and DFDC, SJEDD can still successfully interpret the binary classification results by predicting the authenticity of individual face attributes and regions. For example, on FF++, our detector accurately identifies DF and FS as instances of identity manipulation and F2F and NT as examples of expression forgery.
We additionally compare SJEDD with a baseline detector (i.e., ViT-B [100] trained via binary classification on FFSC) using Grad-CAM visualizations [108], where warmer regions are more important. As shown in Fig. 7, SJEDD provides more human-aligned visualizations by accurately attending to semantically meaningful (and manipulated) face regions.
@@ VI Conclusion and Discussion
We have described a semantics-oriented joint embedding DeepFake detection method called SJEDD, along with a dataset expansion technique. The core of SJEDD lies in the use of joint embedding, which reformulates model parameter splitting in multitask learning to a model capacity allocation problem, allowing for end-to-end bi-level optimization. We replaced the cross-entropy loss with the fidelity loss in bi-level optimization, which additionally boosts detection performance.
The present study addresses DeepFake detection purely from a technical standpoint. Future directions in this field are multifaceted, involving advancements in both technology and policy. From a technological perspective, it is recommended to further explore a multimodal approach that integrates audio, visual, and textual analysis, combined with advanced parameter-efficient fine-tuning techniques like LoRA [109], to improve training efficiency and detection accuracy. Detectors should also be able to operate in real-time on edge devices to support live analysis of streaming content in large-scale media platforms. Another important avenue for advancement is to develop detectors capable of continual learning and adaptation to emerging DeepFake techniques, potentially involving human oversight to maintain detection accuracy and efficacy. Proactive measures such as blockchain technology and digital watermarking to authenticate original content should also be considered. From a policy perspective, it is essential to establish international standards and frameworks for DeepFake detection. This would involve collaboration between governments, tech companies, and research institutions. Laws and regulations should also be enacted to address the creation and distribution of DeepFake content, including penalties for misuse and support for victims.
Acknowledgments
This work was supported in part by the Hong Kong RGC General Research Fund (11220224), the CityU Strategic Research Grants (7005848 and 7005983), and an Industry Gift Fund (9229179).